show that linear combination of eikx and e-ikx is a eigen function of operator d2/dx2

1
Expert's answer
2021-10-06T08:17:05-0400

Let's make the combination:


"f(x) = Ae^{ikx} + Be^{-ikx}"

and apply the operator d2/dx2 to it:


"\dfrac{d^2}{dx^2}[f(x)] = \dfrac{d^2}{dx^2}[Ae^{ikx} + Be^{-ikx}] = \dfrac{d^2}{dx^2}[Ae^{ikx}] + \dfrac{d^2}{dx^2}[ Be^{-ikx}] =\\
=(ik)^2Ae^{ikx} + (-ik)^2Be^{-ikx} = -k^2Ae^{ikx}-k^2Be^{-ikx} =\\
= -k^2(Ae^{ikx} + Be^{-ikx}) = -k^2f(x)"

Thus, obtain:


"\dfrac{d^2}{dx^2}[f(x)] = -k^2f(x)"

By definition, the function "f(x)" is an egenfunction of some operator "A" if "A[f(x)] = \lambda f(x)", which is the case here. Thus, the linear combination of eikx and e-ikx is a eigen function of operator d2/dx2.

Q.E.D.